# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.2.2 Uniform estimates
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{G}^{\gamma,k}+\frac{1}{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\leq\mathcal{H}_{1}^{\gamma,k}\mathcal{G}^{\gamma,k}+\mathcal{H}^{\gamma,k}_{2},$ (3.69)
@@
$\mathcal{H}^{\gamma,k}_{2}(t)$ $=C\gamma^{2}\|\partial_{t}\bm{v}^{k}(t)\|^{2}+\gamma C_{5}\|\bm{v}^{k}(t)\|_{\bm{L}^{\infty}}^{2}\mathcal{H}^{k}_{1}(t)+C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{\gamma,k}}-c_{0}|\mathcal{H}^{k}_{1}(t)$
$\quad+C\big{(}\|\nabla\bm{v}^{k}(t)\|^{2}+\|\nabla\mu^{\gamma,k}(t)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}+\|\nabla(\sigma^{\gamma,k}(t)-\chi\varphi^{\gamma,k}(t))\|^{2}\big{)}$
$\quad+C(\alpha^{2}+\alpha^{4})|\overline{\varphi^{\gamma,k}(t)}-c_{0}|.$
Since $\bm{v}^{k}\in C_{0}^{\infty}((0,+\infty);D(\bm{S}))$, for any $T\in(0,+\infty)$, we infer from (3.39) and (3.42) that both $\mathcal{H}^{k}_{1}$ and $\mathcal{H}^{k}_{2}$ belong to $L^{1}(0,T)$. Applying Gronwall’s lemma, we infer from (3.69) that
$\mathcal{G}^{\gamma,k}(t)\leq\left(\mathcal{G}^{\gamma,k}(0)+\int_{0}^{t}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s\right)\exp\left(\int_{0}^{t}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s\right),\quad\forall\,t\in[0,T],$
with (noticing that $\bm{v}^{k}(0)=\mathbf{0}$)
$\mathcal{G}^{\gamma,k}(0)=\frac{1}{2}\|\nabla\mu^{\gamma,k}(0)\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}(0)\|^{2}+\alpha(\overline{\varphi_{0,k}}-c_{0})\int_{\Omega}\mu^{\gamma,k}(0)\,\mathrm{d}x$
Using (3.67) and (3.68), we arrive at
$\|\nabla\mu^{\gamma,k}(t)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}$
$\quad\leq 4\left(\mathcal{G}^{\gamma,k}(0)+\int_{0}^{t}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s\right)\exp\left(\int_{0}^{t}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s\right)$
$\qquad+4\gamma C_{5}\|\bm{v}^{k}(t)\|_{\bm{L}^{\infty}(\Omega)}^{2}+4C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{\gamma,k}(t)}-c_{0}|,\quad\forall\,t\in[0,T].$ (3.70)
Moreover, integrating (3.69), we obtain
$\int_{0}^{T}\|\nabla\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s$ $\leq 2\mathcal{G}^{\gamma,k}(0)+2\sup_{t\in[0,T]}\mathcal{G}^{\gamma,k}(t)\int_{0}^{T}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s+2\int_{0}^{T}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s$
$\quad+2\gamma C_{5}\|\bm{v}^{k}\|_{L^{\infty}(0,T;\bm{L}^{\infty}(\Omega))}^{2}+4C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{\gamma,k}(T)}-c_{0}|.$ (3.71)
It remains to estimate the initial datum $\mathcal{G}^{\gamma,k}(0)$. Define
$\mu_{0,k}=-\Delta\varphi_{0,k}+\varPsi^{\prime}(\varphi_{0,k})-\chi\sigma_{0,k}+\beta\mathcal{N}(\varphi_{0,k}-\overline{\varphi_{0,k}}).$
By definition, we have
$\|\nabla\mu_{0,k}\|\leq C\big{(}\|\nabla\widetilde{\mu}_{0}\|+\|\varphi_{0}\|_{H^{1}}+\|\sigma_{0}\|_{H^{1}}\big{)},$
where $C>0$ is independent of the parameters $\gamma$ and $k$. Noticing that $\partial_{t}\varphi^{\gamma,k}$, $\mu^{\gamma,k}\in C([0,T];H^{1}(\Omega))$, $\sigma^{\gamma,k}\in C([0,T];H^{2})$, we infer from (3.36b) that $-\Delta\varphi^{\gamma,k}+\varPsi^{\prime}(\varphi^{\gamma,k})\in C([0,T];H^{1}(\Omega))$. Following a similar argument for [30, (3.46)], we can deduce that (using again $\bm{v}^{k}(0)=\mathbf{0}$):
$\|\nabla\mu^{\gamma,k}(0)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(0)\|^{2}$ $=\gamma\alpha^{2}|\overline{\varphi_{0,k}}-c_{0}|^{2}+\int_{\Omega}\nabla\mu^{\gamma,k}(0)\cdot\nabla\mu_{0,k}\,\mathrm{d}x$
$\leq\alpha^{2}|\overline{\varphi_{0,k}}-c_{0}|^{2}+\frac{1}{2}\|\nabla\mu^{\gamma,k}(0)\|^{2}+\frac{1}{2}\|\nabla\mu_{0,k}\|^{2}.$
Applying (3.48), we also find
$\left|\int_{\Omega}\mu^{\gamma,k}(0)\,\mathrm{d}x\right|\leq C\big{(}1+\|\nabla\mu^{\gamma,k}(0)\|+\gamma\|\partial_{t}\varphi^{\gamma,k}(0)\|\big{)}\leq C\|\nabla\mu_{0,k}\|+C,$
which is again independent of $\gamma$ and $k$. Hence, it follows from Young’s inequality that
$\mathcal{G}^{\gamma,k}(0)\leq\|\nabla\mu_{0,k}\|^{2}+C\alpha|\overline{\varphi_{0,k}}-c_{0}|,$ (3.72)
which is uniformly bounded with respective to $\gamma\in(0,1)$ and $k$ (sufficiently large).
3.2.3 Existence and uniqueness

### v2

@@ 3.2.2 Decoupled auxiliary problem for $(\varphi,\mu)$
$\frac{\mathrm{d}}{\mathrm{d}t}\widehat{\mathcal{G}^{\gamma,k}}+\frac{1}{2}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}\leq\mathcal{H}_{1}^{\gamma,k}\widehat{\mathcal{G}^{\gamma,k}}+\mathcal{H}^{\gamma,k}_{2},\$(3.78)
with
$\mathcal{H}^{\gamma,k}_{2}(t)$ $=C(\|\nabla\mu^{\gamma,k}(t)\|^{2}+\gamma^{2}\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2})+C\gamma^{2}\|\partial_{t}\bm{v}^{k}(t)\|^{2}$
$\quad+C(1+\|\nabla\bm{v}^{k}(t)\|^{2})(1+\|\sigma^{k}(t)\|_{L^{6}}^{2})+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2}.$
Here, the generic constant $C>0$ is independent of $\gamma$, $k$ and time. From the construction of $\bm{v}^{k}$, $\sigma^{k}$ and the lower order estimate (3.72), for any given $T\in[1,+\infty)$, we infer that the non-negative functions $\widehat{\mathcal{G}^{\gamma,k}}$, $\mathcal{H}^{\gamma,k}_{1}$, and $\mathcal{H}^{\gamma,k}_{2}$ belong to $L^{1}(0,T)$ with upper bounds possibly depending on $k$, $T$, but not on $\gamma$ due to the fact $\gamma\in(0,1)$.
Let us first consider the estimate on the time interval $[0,1]$. Applying Gronwall’s lemma, we can deduce from (3.78) that
$\sup_{t\in[0,1]}\widehat{\mathcal{G}^{\gamma,k}}(t)\leq\left(\widehat{\mathcal{G}^{\gamma,k}}(0)+\int_{0}^{1}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s\right)\exp\left(\int_{0}^{1}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s\right),$
with the initial datum (noticing that $\bm{v}^{k}(0)=\mathbf{0}$)
$\widehat{\mathcal{G}^{\gamma,k}}(0)$ $=\frac{1}{2}\|\nabla\mu^{\gamma,k}(0)\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}(0)\|^{2}+\alpha(\overline{\varphi_{0,k}}-c_{0})\int_{\Omega}\mu^{\gamma,k}(0)\,\mathrm{d}x$
$\quad+\gamma C_{14}(k,T)+C_{15}.$
As a consequence, we get
$\sup_{t\in[0,1]}\left(\|\nabla\mu^{\gamma,k}(t)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}\right)\leq 4\sup_{t\in[0,1]}\widehat{\mathcal{G}^{\gamma,k}}(t)$
$\qquad\leq 4\left(\widehat{\mathcal{G}^{\gamma,k}}(0)+\int_{0}^{1}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s\right)\exp\left(\int_{0}^{1}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s\right).$ (3.79)
Integrating (3.78) with respect to time, we further obtain
$\int_{0}^{1}\|\nabla\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s$ $\leq 2\widehat{\mathcal{G}^{\gamma,k}}(0)+2\sup_{t\in[0,1]}\widehat{\mathcal{G}^{\gamma,k}}(t)\int_{0}^{1}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s+2\int_{0}^{1}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s.$ (3.80)
Then it is left to estimate the initial datum $\widehat{\mathcal{G}^{\gamma,k}}(0)$. Define
$\widehat{\mu}_{0,k}$ $=-\Delta\varphi_{0,k}+\varPsi^{\prime}(\varphi_{0,k})-\chi\sigma^{k}(0)+\beta\mathcal{N}(\varphi_{0,k}-\overline{\varphi_{0,k}})$
$=\widetilde{\mu}_{0,k}-\theta_{0}\varphi_{0,k}-\chi\sigma^{k}(0)+\beta\mathcal{N}(\varphi_{0,k}-\overline{\varphi_{0,k}}).$
By the definition of $\widetilde{\mu}_{0,k}$ and the construction of $\varphi_{0,k}$, $\sigma^{k}$, we find
$\|\nabla\widehat{\mu}_{0,k}\|\leq C_{16}\big(\|\nabla\widetilde{\mu}_{0}\|+\|\varphi_{0}\|_{H^{1}}+\|\sigma(0)\|_{H^{1}}\big),$ (3.81)
where $C_{16}>0$ is independent of $\gamma$, $k$ and time. Noticing that $\partial_{t}\varphi^{\gamma,k}$, $\mu^{\gamma,k}\in C([0,T];H^{1}(\Omega))$ and $\sigma^{k}\in C([0,T];H^{1}(\Omega))$, we infer from equation (3.66b) that $\Delta\varphi^{\gamma,k}+\varPsi^{\prime}(\varphi^{\gamma,k})\in C([0,T];H^{1}(\Omega))$. Following a similar argument for [34, (3.46)], we can deduce that (using again $\bm{v}^{k}(0)=\mathbf{0}$):
$\|\nabla\mu^{\gamma,k}(0)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(0)\|^{2}$ $=\gamma\alpha^{2}|\Omega||\overline{\varphi_{0,k}}-c_{0}|^{2}+\int_{\Omega}\nabla\mu^{\gamma,k}(0)\cdot\nabla\widehat{\mu}_{0,k}\,\mathrm{d}x$
$\leq 4\alpha^{2}|\Omega|+\frac{1}{2}\|\nabla\mu^{\gamma,k}(0)\|^{2}+\frac{1}{2}\|\nabla\widehat{\mu}_{0,k}\|^{2}.$
From the above estimate, (3.70), $\gamma\in(0,1)$, $\|\varphi_{0,k}\|_{L^{\infty}}\leq 1$ and Young’s inequality, we also obtain
$\left|\int_{\Omega}\mu^{\gamma,k}(0)\,\mathrm{d}x\right|$ $\leq C\big(1+\|\nabla\mu^{\gamma,k}(0)\|+\gamma\|\partial_{t}\varphi^{\gamma,k}(0)\|+\|\sigma^{k}(0)\|_{H^{1}}\big)$
$\leq\frac{1}{2}\|\nabla\widehat{\mu}_{0,k}\|^{2}+C(1+\|\sigma(0)\|_{H^{1}}),$
which is again independent of $\gamma$ and $k$ (cf. (3.81)). In summary, we get
$\widehat{\mathcal{G}^{\gamma,k}}(0)\leq\|\nabla\widehat{\mu}_{0,k}\|^{2}+8\alpha^{2}|\Omega|+2\gamma C_{14}(k,T)+2C_{15},$ (3.82)
which is uniformly bounded with respect to $\gamma\in(0,1)$.
Next, we consider the estimate on $[1,T]$. An application of the uniform Gronwall inequality (see, e.g., [52, Chapter III, Lemma 1.1]) to (3.78) yields
$\widehat{\mathcal{G}^{\gamma,k}}(t)$ $\leq\left(\sup_{r\in[0,T-1]}\int_{r}^{r+1}\widehat{\mathcal{G}^{\gamma,k}}(s)\,\mathrm{d}s+\sup_{r\in[0,T-1]}\int_{r}^{r+1}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s\right)$
$\quad\ \ \times\exp\left(\sup_{r\in[0,T-1]}\int_{r}^{r+1}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s\right),\quad\forall\,t\in[1,T].$ (3.83)
Then integrating (3.78) with respect to time, we get
$\int_{t}^{t+1}\|\nabla\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s$ $\leq 2\widehat{\mathcal{G}^{\gamma,k}}(t)+2\sup_{r\in[t,t+1]}\widehat{\mathcal{G}^{\gamma,k}}(r)\int_{t}^{t+1}\mathcal{H}^{\gamma,k}_{1}(s)\,\mathrm{d}s$
$\quad+2\int_{t}^{t+1}\mathcal{H}^{\gamma,k}_{2}(s)\,\mathrm{d}s,\quad\forall\,t\in[0,T-1].$ (3.84)
Combining (3.72), (3.79)–(3.84), we arrive at higher-order estimates for $(\varphi^{\gamma,k},\mu^{\gamma,k})$ on $[0,T]$ (note that $T\geq 1$ is arbitrary but fixed), which are uniform with respect to the parameter $\gamma\in(0,1)$. Recalling (3.74), we further obtain
$\sup_{t\in[0,T]}\|\varphi^{\gamma,k}(t)\|_{H^{2}}+\sup_{t\in[0,T-1]}\int_{t}^{t+1}\left(\|\varphi^{\gamma,k}(s)\|_{W^{2,6}}^{2}+\|\varPsi_{0}^{\prime}(\varphi^{\gamma,k}(s))\|_{L^{6}}^{2}\right)\,\mathrm{d}s\leq C.$ (3.85)
The vanishing viscosity limit as $\gamma\to 0^{+}$. Thanks to the estimates (3.68), (3.72), (3.74) and (3.79)–(3.85), we can pass to the limit as $\gamma\to 0^{+}$ (keeping $k$, $T$ fixed) such that there exists a convergent subsequence $\{(\varphi^{\gamma,k},\mu^{\gamma,k})\}$ (not relabeled for simplicity) with the limit denoted by $(\varphi^{k},\mu^{k})$. The procedure follows a standard compactness argument (see e.g., [3, 25]) and its details are omitted here. The limit triple $(\varphi^{k},\mu^{k})$ satisfies
$\varphi^{k}\in L^{\infty}(0,T;H^{2}(\Omega))\cap L^{2}(0,T;W^{2,6}(\Omega))\cap H^{1}(0,T;H^{1}(\Omega)),$
$\mu^{k}\in L^{\infty}(0,T;H^{1}(\Omega)),$
$\varphi^{k}\in L^{\infty}(\Omega\times(0,T))\ \ \textrm{with}\ \ |\varphi^{k}(x,t)|<1\quad\textrm{a.e.\ in}\ \Omega\times(0,T),$
for any $T>0$. Noticing that $(\varphi^{k},\mu^{k})$ satisfies
$(\partial_{t}\varphi^{k},\xi)+(\bm{v}^{k}\cdot\nabla\varphi^{k},\xi)=-(\nabla\mu^{k},\nabla\xi)-\alpha\big((\overline{\varphi^{k}}-c_{0}),\xi\big),\quad\forall\,\xi\in H^{1}(\Omega),$
almost everywhere in $(0,T)$, and
$\|\bm{v}^{k}\cdot\nabla\varphi^{k}\|$ $\leq C\|\bm{v}^{k}\|_{\bm{L}^{6}}\|\varphi^{k}\|_{W^{1,3}}\leq C\|\bm{v}^{k}\|_{\bm{H}^{1}}\|\varphi^{k}\|_{H^{2}},$
we can infer from the elliptic regularity theory that
$\mu^{k}\in L^{2}(0,T;H^{2}(\Omega)).$
With the above estimates, it is straightforward to check that the strong solution $(\varphi^{k},\mu^{k})$ is unique. In addition, it solves the following system
$\begin{cases}&\partial_{t}\varphi^{k}+\bm{v}^{k}\cdot\nabla\varphi^{k}=\Delta\mu^{k}-\alpha(\overline{\varphi^{k}}-c_{0}),\\ &\mu^{k}=-\Delta\varphi^{k}+\varPsi^{\prime}(\varphi^{k})-\chi\sigma^{k}+\beta\mathcal{N}(\varphi^{k}-\overline{\varphi^{k}}),\end{cases}$ (3.86)
almost everywhere in $\Omega\times(0,T)$, and fulfills the boundary and initial conditions:
$\begin{cases}{\partial}_{\bm{n}}\varphi^{k}={\partial}_{\bm{n}}\mu^{k}=0,\quad\textrm{a.e. on}\ \partial\Omega\times(0,T),\\ \varphi^{k}|_{t=0}=\varphi_{0,k}(x),\quad\ \ \,\textrm{a.e. in}\ \Omega.\end{cases}$ (3.87)
Since $T>0$ is arbitrary, the solution $(\varphi^{k},\mu^{k})$ can be uniquely extended to $0,+\infty)$. However, at this stage, we only have estimates in an arbitrary but fixed interval $[0,T]$.
Passage to the limit as $k\to+\infty$. For every solution $(\varphi^{k},\mu^{k})$, we can repeat the previous argument without the viscous term (i.e., $\gamma=0$) to show that lower-order estimates corresponding to (3.68), (3.72), (3.74) are still valid, being uniform with respect to the parameter $k$ and the final time $T$.
In what follows, we derive higher-order estimates for $(\varphi^{k},\mu^{k})$. We shall work on the convergent subsequence $\big\{(\varphi^{\gamma,k},\mu^{\gamma,k})\big\}$ (not relabeled for simplicity) with the corresponding limit $(\varphi^{k},\mu^{k})$. All convergence results mentioned in the following should be understood in the sense of a subsequence.
Integrating (3.69) with respect to time, the approximate solution $(\varphi^{\gamma,k},\mu^{\gamma,k})$ satisfies the following energy identity:
$\frac{1}{2}\|\nabla\varphi^{\gamma,k}(t_{2})\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k}(t_{2}))\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}(t_{2})-\overline{\varphi^{\gamma,k}(t_{2})})\|^{2}$
$\qquad+\int_{t_{1}}^{t_{2}}(\|\nabla\mu^{\gamma,k}(s)\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}(s)\|^{2})\,\mathrm{d}s$
$\quad=\frac{1}{2}\|\nabla\varphi^{\gamma,k}(t_{1})\|^{2}+\int_{\Omega}\varPsi(\varphi^{\gamma,k}(t_{1}))\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{\gamma,k}(t_{1})-\overline{\varphi^{\gamma,k}(t_{1})})\|^{2}$
$\qquad-\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}(\bm{v}^{k}(s)\cdot\nabla\varphi^{\gamma,k}(s))(\mu^{\gamma,k}(s)+\chi\sigma^{k}(s))\,\mathrm{d}x\mathrm{d}s-\chi\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\nabla\mu^{\gamma,k}(s)\cdot\nabla\sigma^{k}(s)\,\mathrm{d}x\mathrm{d}s$
$\qquad-\int_{t_{1}}^{t_{2}}\alpha(\overline{\varphi^{\gamma,k}(s)}-c_{0})\int_{\Omega}(\mu^{\gamma,k}(s)+\chi\sigma^{k}(s))\,\mathrm{d}x\mathrm{d}s,$
for any $0\leq t_{1}\leq t_{2}\leq T$. Due to the regularity of $(\varphi^{k},\mu^{k})$, we can check that it also satisfies an energy identity:
$\frac{1}{2}\|\nabla\varphi^{k}(t_{2})\|^{2}+\int_{\Omega}\varPsi(\varphi^{k}(t_{2}))\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{k}(t_{2})-\overline{\varphi^{k}(t_{2})})\|^{2}+\int_{t_{1}}^{t_{2}}\|\nabla\mu^{k}(s)\|^{2}\,\mathrm{d}s$
$\quad=\frac{1}{2}\|\nabla\varphi^{k}(t_{1})\|^{2}+\int_{\Omega}\varPsi(\varphi^{k}(t_{1}))\,\mathrm{d}x+\frac{\beta}{2}\|\nabla\mathcal{N}(\varphi^{k}(t_{1})-\overline{\varphi^{k}(t_{1})})\|^{2}$
$\qquad-\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}(\bm{v}^{k}(s)\cdot\nabla\varphi^{k}(s))(\mu^{k}(s)+\chi\sigma^{k}(s))\,\mathrm{d}x\mathrm{d}s-\chi\int_{t_{1}}^{t_{2}}\!\!\int_{\Omega}\nabla\mu^{k}(s)\cdot\nabla\sigma^{k}(s)\,\mathrm{d}x\mathrm{d}s$
$\qquad-\int_{t_{1}}^{t_{2}}\alpha(\overline{\varphi^{k}(s)}-c_{0})\int_{\Omega}(\mu^{k}(s)+\chi\sigma^{k}(s))\,\mathrm{d}x\mathrm{d}s,$
for any $0\leq t_{1}\leq t_{2}\leq T$. Using (3.68), (3.80), (3.81), (3.82), (3.84) and the Poincaré–Wirtinger inequality, we observe that
$\int_{t_{1}}^{t_{2}}\gamma\|\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s$ $\leq\gamma C\int_{t_{1}}^{t_{2}}\|\nabla\partial_{t}\varphi^{\gamma,k}(s)\|^{2}\,\mathrm{d}s+\gamma C|\Omega|\int_{t_{1}}^{t_{2}}|\overline{\partial_{t}\varphi^{\gamma,k}(s)}|^{2}\,\mathrm{d}s$
$\to 0,\quad\text{as}\ \gamma\to 0^{+}.$
By the argument in [3, Section 2], comparing the two energy identities above and taking the limit as $\gamma\to 0^{+}$, we obtain
$\lim_{\gamma\to 0^{+}}\int_{t_{1}}^{t_{2}}\|\nabla\mu^{\gamma,k}(s)\|^{2}\,\mathrm{d}s=\int_{t_{1}}^{t_{2}}\|\nabla\mu^{k}(s)\|^{2}\,\mathrm{d}s,$
for any $0\leq t_{1}\leq t_{2}\leq T$. With this observation, by the lower semi-continuity of norms with respect to the weak convergence, we can pass to the limit as $\gamma\to 0^{+}$ in the inequalities (3.79), (3.80) to get (with the aid of (3.81), (3.82))
$\sup_{t\in[0,1]}\|\nabla\mu^{k}(t)\|^{2}\leq 4\left(C_{17}+\int_{0}^{1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s\right)\exp\left(C\int_{0}^{1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right),$
and
$\int_{0}^{1}\|\nabla\partial_{t}\varphi^{k}(s)\|^{2}\,\mathrm{d}s$ $\leq 2C_{17}+2C\left(C_{17}+\int_{0}^{1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s\right)\int_{0}^{1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s$
$\qquad\ \times\exp\left(C\int_{0}^{1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right)+2\int_{0}^{1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s,$
@@
$C_{17}$ $=C_{16}\big(\|\nabla\widetilde{\mu}_{0}\|+\|\varphi_{0}\|_{H^{1}}+\|\sigma(0)\|_{H^{1}}\big)+8\alpha^{2}|\Omega|+2C_{15},$
$\mathcal{H}^{k}_{2}(t)$ $=C\|\nabla\mu^{k}(t)\|^{2}+C(1+\|\nabla\bm{v}^{k}(t)\|^{2})(1+\|\sigma^{k}(t)\|_{L^{6}}^{2})+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2}.$
Similarly, we can pass to the limit as $\gamma\to 0^{+}$ in the inequalities (3.83), (3.84) to obtain (with the aid of (3.81), (3.82))
$\sup_{t\in[1,T]}\|\nabla\mu^{k}(t)\|^{2}$ $\leq 4\left(\sup_{r\in[0,T-1]}A(r)+\sup_{r\in[0,T-1]}\int_{r}^{r+1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s\right)$
$\quad\ \ \ \times\exp\left(C\sup_{r\in[0,T-1]}\int_{r}^{r+1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right),$
and
$\sup_{t\in[0,T-1]}\int_{t}^{t+1}\|\nabla\partial_{t}\varphi^{k}(s)\|^{2}\,\mathrm{d}s$
$\quad\leq 2\left(C_{17}+\sup_{r\in[0,T-1]}A(r)+\sup_{r\in[0,T-1]}\int_{r}^{r+1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s\right)\left(1+C\int_{t}^{t+1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right)$
$\qquad\ \times\exp\left(C\sup_{r\in[0,T-1]}\int_{r}^{r+1}\|\nabla\bm{v}^{k}(s)\|^{2}\,\mathrm{d}s\right)+2\int_{t}^{t+1}\mathcal{H}^{k}_{2}(s)\,\mathrm{d}s,$
where
$A(r)=\int_{r}^{r+1}\|\nabla\mu^{k}(s)\|^{2}\,\mathrm{d}s+2C_{15},\quad r\in[0,T-1].$
Hence, we have derived estimates for $\mu^{k}$, $\partial_{t}\varphi^{k}$ on $[0,T]$, which are independent of $k$ and $T$.
Collecting the above estimates, we can conclude
$\|\mu^{k}\|_{L^{\infty}(0,+\infty;H^{1}(\Omega))}+\|\partial_{t}\varphi^{k}\|_{L^{2}_{\mathrm{uloc}}(0,+\infty;H^{1}(\Omega))}$
$\quad+\|\varphi^{k}\|_{L^{\infty}(0,+\infty;W^{2,6}(\Omega))}+\|\varPsi^{\prime}(\varphi^{k})\|_{L^{\infty}(0,+\infty;L^{6}(\Omega))}\leq C.$ (3.88)
Noticing that
$\|\bm{v}^{k}\cdot\nabla\varphi^{k}\|_{H^{1}}$ $\leq C(\|\bm{v}^{k}\|_{\bm{H}^{1}}\|\nabla\varphi^{k}\|_{\bm{L}^{\infty}}+\|\bm{v}^{k}\|_{\bm{L}^{6}}\|\varphi^{k}\|_{W^{2,3}})$
$\leq C\|\bm{v}^{k}\|_{\bm{H}^{1}}\|\varphi^{k}\|_{W^{2,6}},$
from (3.88) and the elliptic regularity theory applied to the first equation in (3.86), we also get
$\|\mu^{k}\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);H^{3}(\Omega))}\leq C.$ (3.89)
The positive constant $C$ appearing in (3.88), (3.89) depends on $\max_{r\in[-1,1]}|\varPsi(r)|$, $\|\widetilde{\mu}_{0}\|_{H^{1}}$, $\|\varphi_{0}\|_{H^{1}}$, $\overline{\varphi_{0}}$, $\|\sigma(0)\|_{H^{1}}$, the coefficients of the system, $\Omega$ and norms of the given data $\|\bm{v}\|_{L^{2}(0,+\infty;\bm{H}_{0,\mathrm{div}}^{1}(\Omega))}$, $\|\sigma\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);H^{1}(\Omega))}$, $\|\sigma\|_{L^{\infty}(0,+\infty;L^{6}(\Omega))}$ and $\|\partial_{t}\sigma\|_{L^{2}_{\mathrm{uloc}}([0,+\infty);(H^{1}(\Omega))^{\prime})}$, but it is independent of $k$ and $T$.
With the above uniform estimates, we can apply the standard compactness argument and pass to the limit as $k\to+\infty$. There exists a convergent subsequence $\big\{(\varphi^{k},\mu^{k})\}$ whose limit is denoted by $(\varphi,\mu)$. It is straightforward to check that $(\varphi,\mu)$ is a solution to the original problem (3.55) on $0,+\infty)$ with the corresponding regularity properties as stated in Proposition 3.2-(2).
By comparison in (3.1a), we find
Thus, if in addition, $\bm{v}\in L^{\infty}(0,+\infty;\bm{L}^{2}_{0,\mathrm{div}}(\Omega))$, then $\partial_{t}\varphi\in L^{\infty}(0,+\infty;(H^{1}(\Omega))^{\prime})$.
The proof of Proposition 3.2-(2) is complete. ∎
3.2.3 Proof of Theorem 3.2
We are ready to prove Theorem 3.2.